# The discrete nonlinear Schrödinger equation with linear gain and nonlinear loss: the infinite lattice with nonzero boundary conditions and its finite dimensional approximations

### v1

@@ VI Conclusions
In this work, we studied the dynamics of the Discrete Nonlinear Schrödinger Equation incorporating linear gain and nonlinear loss effects, a significant model in various physical contexts. In the case of the infinite lattice we considered the problem with nonzero boundary conditions proving that solutions exist if and only if the supporting background has a prescribed critical value $A_{*}$ defined by the gain and loss strengths. We argued that this property cannot be captured by finite lattice approximations, as those defined by the system supplemented with periodic boundary conditions, which are essential for the numerical simulations of the problem. For the periodic lattice, which is of physical and mathematical significance itself, we proved that the dynamics are dominated by the convergence to a global attractor which is the unique plane wave of constant amplitude, namely $A_{*}$, with a novel argument which studies not only the limits of the amplitudes of the initial conditions but also their frequencies and spectrum. The convergence arguments accompanied by a modulation instability analysis justify that modulation instability, the fundamental mechanism for the emergence of localized structures can be only transient in the case of the periodic lattice. The dynamics of localized initial data are investigated numerically: in the case where $A=A_{*}$ which is the only relevant in order to approximate the infinite lattice, we identify the persistence of the major characteristics of the modulation instability patterns manifested in the dynamics of integrable NLS systems, as the Ablowitz-Ladik lattice: the oscillatory wedge shaped sector and the two outer plane wave regions of the undisturbed background of amplitude $A_{*}$, which is globally asymptotically stable in the case of the finite lattice. In the case $A\neq A_{*}$ which is only relevant for the finite lattice, the modulation instability pattern still persists, this time on a varying background transiently, prior the convergence of the dynamics to the plane wave attractor. At the early stages of the dynamics we identified the emergence of rogue wave events which are remarkably proximal to the analytical discrete Peregrine rogue wave solution of the Ablowitz-Ladik lattice. This proximity is justified analytically by suitable estimates for the distance between the solutions of the DNLS system and the Ablowitz-Ladik lattice.

### v3

@@ 2. We start by seeking solutions of the form
V.5 Concluding remarks for the boundary conditions
  1. 1.
    Generalized boundary conditions at infinity: existence and uniqueness. Motivated again by [13] and [21], instead of the boundary conditions (1.2), one may seek to consider more general nonzero boundary conditions at infinity, of the form
    $\lim_{|n|\rightarrow\infty}u_{n}(t)=\lim_{|n|\rightarrow\infty}e^{iG^{2}t}\zeta_{n},$ (5.33)
    where $G\in\mathbb{R}$ and
    $\zeta_{n}=\left\{\begin{array}[]{ll}\zeta_{-},&n<0,\\ \zeta_{+},&n\geq 0,\end{array}\right.\;\;\;\mbox{$\zeta_{\pm}\in\mathbb{C}$ with $|\zeta_{\pm}|=\zeta$.}$ (5.34)
    Using first the change of variables $u_{n}(t)=\psi_{n}(t)\exp(\mathrm{i}G^{2}t)$, and then, the change of variables $U_{n}={\psi}_{n}-\zeta_{n}$, we derive that the modified system for $U_{n}$ reads now as
    $\mathrm{i}\dot{U}_{n}+\Delta_{d}U_{n}+\Delta_{d}\zeta_{n}-G^{2}(U_{n}+\zeta_{n})$ $+|U_{n}+\zeta_{n}|^{2}(U_{n}+\zeta_{n})$
    $=\mathrm{i}\gamma(U_{n}+\zeta_{n})+\mathrm{i}\delta|U_{n}+\zeta_{n}|^{2}(U_{n}+\zeta_{n}),\;\;$ (5.35)
    and the boundary conditions for $U_{n}$ are again zero at infinity, that is, the solution of (1) must satisfy (2.10). Note that the term $\Delta_{d}\zeta_{n}\in\ell^{2}$ by the definition of $\zeta_{n}$ in (5.34). The operators $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are defined in this case by
    $\mathcal{F}_{1}(U_{n})$ $=-G^{2}(U_{n}+\zeta_{n})+|U_{n}+\zeta_{n}|^{2}(U_{n}+\zeta_{n}),$
    $\mathcal{F}_{2}(U_{n})$ $=\mathrm{i}\gamma(U_{n}+\zeta_{n})+\mathrm{i}\delta|U_{n}+\zeta_{n}|^{2}(U_{n}+\zeta_{n}).$
    The operator $\mathcal{F}_{1}$ can be written as
    $\mathcal{F}_{1}(U_{n})=-G^{2}U_{n}-G^{2}\zeta_{n}+|\zeta_{n}|^{2}\zeta_{n}+\left\{|U_{n}|^{2}U_{n}+2(|U_{n}|^{2}\zeta_{n}+U_{n}|\zeta_{n}|^{2})+U_{n}^{2}\overline{\zeta}_{n}+\overline{U}_{n}\zeta_{n}^{2}\right\},$
    while the operator $\mathcal{F}_{2}$, as
    $\mathcal{F}_{2}(U_{n})=\mathrm{i}\gamma U_{n}+\mathrm{i}\gamma\zeta_{n}+\mathrm{i}\delta|\zeta_{n}|^{2}\zeta_{n}+\mathrm{i}\delta\left\{|U_{n}|^{2}U_{n}+2(|U_{n}|^{2}\zeta_{n}+U_{n}|\zeta_{n}|^{2})+U_{n}^{2}\overline{\zeta}_{n}+\overline{U}_{n}\zeta_{n}^{2}\right\}.$
    Carrying out the same analysis as for the proof of Theorem II.1 for the operators $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, we have that a unique solution $U\in C^{1}([0,T],\ell^{p})$ of the modified equation (1) exists, if and only if
    $-G^{2}\zeta_{n}+|\zeta_{n}|^{2}\zeta_{n}=0\quad\text{and}\quad\mathrm{i}\gamma\zeta_{n}+\mathrm{i}\delta|\zeta_{n}|^{2}\zeta_{n}=0.$
    Therefore, we conclude with the following corollary.
    Corollary V.1.
    Consider the DNLS system (1.1) with $\gamma>0,\delta<0$ in the infinite lattice supplemented with the nonzero boundary conditions (5.33)-(5.34). We also assume that the initial condition satisfies the boundary conditions (5.33)(5.34). Then the system has a solution if and only if $G^{2}=\zeta^{2}$ and $\zeta=A_{*}$. The solution is unique.
  2. 2.
    Dirichlet boundary conditions. (a) As proven in Theorem II.1, Dirichlet boundary conditions are relevant for approximating problem (1.1)-(1.2) if and only if one uses the equivalent modified DNLS equation (II) supplemented with zero Dirichlet boundary conditions. This approach approximates the zero boundary conditions at infinity (2.10) for (II). Repeating the numerical studies of sections IV and V with this approximation and using the change of variables $u_{n}(t)=(U_{n}(t)+A)\exp(\mathrm{i}A^{2}t)$ produces exactly the same results as those reported in these sections for the DNLS (1.1) supplemented with periodic boundary conditions. (b) Clearly, Dirichlet boundary conditions for the original, unmodified DNLS (1.1) are not relevant for the finite lattice approximation of the non-zero boundary conditions at infinity (1.2).
  3. 3.
    Generalized boundary conditions: asymptotic behavior. Corollary V.1 generalizes Theorem II.1 to the case of more general nonzero boundary conditions at infinity (5.33)-(5.34). Then, one can speculate that the results on the structure of the attractor of section III and the numerical results of sections IV and V can be of similar nature as for the simplest case of boundary conditions (1.2). However, this speculation, to be verified, needs theoretical and numerical analysis either (i) on the modified DNLS (1) supplemented with Dirichlet boundary conditions or (ii) on the unmodified DNLS (1.1) supplemented, this time, with the relevant non-zero Dirichlet boundary conditions. Such an analysis may have similarities but also differences compared with the one of the boundary conditions (1.2) and may deserve independent attention to be considered elsewhere. In this regard, even more general boundary conditions can be considered with a varying $\zeta_{n}$, such that $\lim_{n\rightarrow\infty}|\zeta_{n}|=\zeta$ (a discretized version of the boundary conditions of [28, Section 3, pg. 127]).
@@ VI Conclusions
In this work, we studied the dynamics of the Discrete Nonlinear Schrödinger Equation incorporating linear gain and nonlinear loss effects, a significant model in various physical contexts. In the case of the infinite lattice we considered the problem with nonzero boundary conditions proving that solutions exist if and only if the supporting background has a prescribed critical value $A_{*}$ defined by the gain and loss strengths. We argued that this property cannot be captured by finite lattice approximations, as those defined by the system supplemented with periodic boundary conditions, which are essential for the numerical simulations of the problem. For the periodic lattice, which is of physical and mathematical significance itself, we proved that the dynamics are dominated by the convergence to a global attractor which is the unique plane wave of constant amplitude, namely $A_{*}$, with a novel argument which studies not only the limits of the amplitudes of the initial conditions but also their frequencies and spectrum. The convergence arguments accompanied by the analysis for modulation instability justify that this fundamental mechanism for the emergence of localized structures can be only transient in the case of the periodic lattice. The dynamics of localized initial data are investigated numerically: in the case where $A=A_{*}$ which is the only relevant in order to approximate the infinite lattice, we identify the persistence of the major characteristics of the modulation instability patterns manifested in the dynamics of integrable NLS systems, as the Ablowitz-Ladik lattice: the oscillatory wedge shaped sector and the two outer plane wave regions of the undisturbed background of amplitude $A_{*}$, which is globally asymptotically stable in the case of the finite lattice. In the case $A\neq A_{*}$ which is only relevant for the finite lattice, the instability pattern still persists, this time on a varying background transiently, prior the convergence of the dynamics to the plane wave attractor. We may conclude that the analysis of the infinite lattice “leaves its mark” on the dynamics of the finite-dimensional one, particularly regarding the manifestation of the major characteristics of the universal behavior in the nonlinear stage of modulation instability.
At the early stages of the dynamics we identified the emergence of rogue wave events which are remarkably proximal to the analytical discrete Peregrine rogue wave solution of the Ablowitz-Ladik lattice. This proximity is justified analytically by suitable estimates for the distance between the solutions of the DNLS system and the Ablowitz-Ladik lattice. This proximity, in light of the theoretical analysis for the infinite lattice, may pave the way for new and interesting investigations for the potential construction of rogue wave-like solutions for the dissipative DNLS via fixed-point iterations, using the numerical methods outlined in [18, 19, 20].